ING Mechanical Engineering

Minutes of the Mechanical Group Meeting


 

Minutes of the Mechanical Group Meeting

Date 18 January 2001

 

Attendance

K Dee, P Morrall, M V D Hoeven, M Blanken, P Jolley, S Rodriguez, J Carlos (Part Time), F Concepcion (Part Time), D Gonzalez (Part Time)

 

No

Details

Action

     

1

Minutes of the previous meeting 9/11/00

 
 

These were accepted as true

 
     
     

2

Action items

 

2.1

NAOMI Trolley

 
 

Perspex for cover was not machinable. PSM and SRB to source different grade

PSM, SRB

 

Requires brakes - Done

 
 

Requires a dedicated lifting frame - Done

 
 

Following the recent instrument change, the procedures need updating - no progress

JOLLEY

 

Removal of INGRID from NAOMI is difficult. New lifting arrangement for INGRID to be defined - no progress SRB to review and modify parts

KMD, SRB

     

2.2

Cable Wrap

 
 

Document containing cable descriptions and positions to be updated. CMWJ to be contacted - no progress

KMD

     

2.3

Second Optical GHRIL Table

 
 

Following recent problems with the old optical table with SCAM, the provision of a more stable table to be reviewed. New table ordered delivery 6-8 weeks.

KMD

     

2.4

INGRID

 
 

INGRID bracket needs reviewing to include easier initial alignment and fine adjustments. Design completed SRB to source manufacture

SRB

     

2.5

GRHIL

 
 

GRHIL roof safe working to be reviewed. Harnesses to be used - paper note to be distributed. Harness now available, awaiting paper note.

KMD

 

GRHIL roof to have dedicated strops - to be ordered Done

PSM

 

Roof access to be assessed and modifications made. Risk assessment required.

KMD, SRB

     

2.6

Snow cleaning

 
 

Need a regime for snow cleaning. MFB to ensure its scheduled and completed with the aid of the telescope manager and ops team leader. Ongoing

MFB

 

Note to be issued clarifying the importance of snow cleaning and the new procedures for doing so. No progress

MFB/ KMD

     
     

3

ING Overview

 
 

ING funding for the next 5 years have been secured. The board has agreed upon the business plan.

 
     
     

4

Project Reports

 

4.1

INT Counterbalance

 
 

Requires some modifications to allow working in all orientations. New gearboxes to be fitted and protective covers required. Gearboxes ordered, covers might need mods.

PSM

 

PSM to explain instrument change procedure to SRB. Done

PSM

 

PSM to replace centre screws Done

PSM

 

PSM to install protective rubber bounce pads – PSM sourcing suitable rubber

PSM

 

Instrument change document needs updating. MVDH to do and give copy to EJM

MVDH

     

4.2

Safety

 
 

Jolley to convert and publish test certificates on the intranet. September’s complete, no progress on April’s

JOLLEY

 

PSM to liase with AKC to update the lifting document on the intranet. The new document to include authorised lifting personnel, the equipment that they can use and there training requirements/schedules. Done

PSM

 

Lifting Store needs sorting out. PSM and JC to implement a routine for everyone to follow. No progress

PSM/JC

     

4.4

Red CCD's

 
 

David's manufacturing these items expected complete mid December. To be handed over mid/end January

David

     

4.5

Small Fibres

 
 

Prism alignment going well. Jig to be finished. Test to be performed. Done

MVDH/ MFB

 

New fingers/slit unit to be ordered. Done

JOLLEY

 

MFB/JOLLEY to arrange coating of the microlenses. In Progress

MFB/ JOLLEY

 

Transmission tests to be finalised No progress

MFB

 

RAW to detail new field plate.

 
 

Program suffering slippage due to late delivery of items. May/June off-telescope testing, end July commissioning

 
     

4.6

OASIS

 
 

This is a new project to the group, so JOLLEY gave a brief explanation of what it is and what we need to provide.

 
     

4.7

Other Projects

 
 

PN Spectrograph - due next semester

 
 

LIRIS - due soon

 
 

Auxport Camera - no resources available. Now called WAPCAM – has low priority to group

 
 

Long Camera - waiting for optical design and resources

 
 

Near Earth Object (NEO) - close to getting funding for JKT

 
 

Ultracam – Vik Dhillon’s 3 colour camera

 
     
     

5

Maintenance

 

5.1

Oil system at the WHT needs inspecting and possible cleaning. No progress

SRB/ KMD

     

5.2

For information only: KMD stated that care should be taken when working around the WHT bull ring. The microswitch has three positions and it is possible to drive the telescope passed its end stops. IF YOU THINK THIS HAS BEEN KNOCKED/MOVED THEN CHECK BEFORE RELEASING TELESCOPE. Done

 
     

5.3

Following the recent request the WHT secondary mirror alignment was checked and found to be acceptable. Done

 
     

5.4

WHT Dome bogie replacement. SRB/PSM to get price

SRB, PSM

     
     

6

Training

 

6.1

All to actively look for relevant training courses and present to SRB and KMD.

ALL

     

6.2

David to attend edgeCAM course

DAVID

     
     

7

Administration

 
 

No Report

 
     
     

8

A.O.B

 
     

8.1

Mechanical Design Support

 
 

Kevin to source extra short term design support. Done

KMD

     

8.2

Requests for any work that is deemed to be an enhancement must come from the EMM and not individuals. Work that is not an enhancement but necessary for the working of the telescope to be carried out as normal. Done

ALL

     

8.3

Fire risk assessment from Doug highlighted risk due to the storage of excessive quantities of oil in the garage. SRB and PSM to resolve with FC

SRB, PSM, FC

     

8.4

Staff are reminded that this is a no smoking site. KMD will pay particular attention to the mechanical workshop and garage. Staff caught smoking will be disciplined accordingly.

ALL

     

8.5

Staff are reminded that they should inform a third party if they are operating machinery. Lone working is not allowed on machine tools.

ALL

     

8.6

FC needs a computer in the garage. Due to the purchase of a new machine, old machines will be passed down releasing one for FC

 
     

9

Next Meeting

 
 

8th March

 

ING Logo
Last Updated: 02/04/2003
Maarten Blanken jolley@ing.iac.es